If the length of a side of a square is increased by 3 cm, the perimeter of the new square is 40 cm more than twice the length of the side of the original square. Find the length of the side of the original square.
Answer by FrankM(1040)

Side is X
Side increased by 3 is X+3
The perimeter, 4(X+3)=40+2X, now solve for X
4X+12=2X+40
2X=28
X=14
If increased to 17, the perimeter, 68, is 40 more than 28. Done.